# Online Multi-modal Root Cause Analysis

### v1

@@ 3.5 Network Propagation-based Root Cause Identification
The propagation of malfunction effects from the root cause to adjacent entities implies that the immediate neighbors of system KPIs may not necessarily be the root causes themselves. To identify the root cause, we initially derive the transition probability matrix based on the causal graph $\mathcal{G}$ and then utilize a random walk with restart method (Tong et al., 2006) to simulate the spread patterns of malfunctions as follows:
$\bm{P_{ij}}=\frac{\beta\mathcal{A}_{j,i}}{\sum_{k=1}^{n}\mathcal{A}_{k,i}}$ (19)
The transition probability matrix $\bm{P}$ is the normalized adjacency matrix signified by the coefficient $\beta\in[0,1]$. During the visiting exploration process, we may restart from the KPI node to revisit other system entities with the probability $c\in[0,1]$. The equation for the random walk with restart is formulated by:
$\bm{r}_{t+1}=(1-c)\bm{P}\bm{r}_{t}+c\bm{r}_{0}$ (20)
where $\bm{r}_{t}$ represents the jumping probability at the $t^{\textrm{th}}$ step, $\bm{r}_{0}$ denotes the initial starting probability, and $c\in[0,1]$ stands for the restart probability. Upon convergence of the jumping probability $\bm{r}_{t}$, the probability scores of the nodes are employed to rank the system entities and the top $k$ entities are selected as the most probable root causes for system failure.
Stopping Criterion. As the number of new data batches increases, the identified causal structure and its associated root cause list may gradually converge. To prevent unnecessary consumption of computing resources, we employ them as indicators for automatic termination of the online RCA process. We use the rank-biased overlap metric (RBO) (Webber et al., 2010) to measure the similarity between two root cause lists, effectively capturing the evolving trend of root cause rankings. Given the rank lists from the previous and current batches, denoted as $R_{t-1}$ and $R_{t}$ respectively, we quantify the similarity between these lists as follows:
$\gamma$ $=\text{RBO}(R_{t-1},R_{t})$ (21)
@@
where $\gamma\in[0,1]$. A higher value of $\gamma$ indicates a greater similarity between the two root cause lists. The online RCA process is terminated when the similarity score $s$ surpasses a predefined threshold.
4 Experiments
@@
4.1 Experimental Setup
Datasets. We evaluate the performance of OCEAN using three public real-world datasets: (1) Product Review***https://lemma-rca.github.io/docs/data.html (Zheng et al., 2024b): A microservice system dedicated to online product reviews, encompassing $234$ pods deployed across $6$ cloud servers. It recorded four system faults between May 2021 and December 2021. (2) Online Boutiquehttps://github.com/IntelligentDDS/Nezha (Yu et al., 2023): A microservice system designed for e-commerce, including five system faults. (3) Train Ticket (Yu et al., 2023): A microservice system for railway ticketing services, also with five system faults. All three datasets contain two modalities: system metrics and system logs.
Evaluation Metrics. We choose four widely-used metrics (Wang et al., 2023c; Meng et al., 2020): (1) Precision@K (PR@K): Measures the accuracy of the top-K predicted root causes. (2) Mean Average Precision@K (MAP@K): Evaluates the overall accuracy of the top-K predicted causes. (3) Mean Reciprocal Rank (MRR): Assesses the ranking capability of the models. (4) Time: Measures the training time (in seconds) for each batch of data. Details on the first three metrics are provided in Appendix D and we provide the time complexity analysis in Appendix B.
Baselines. We compare OCEAN with seven causal discovery based RCA methods: (1) PC (Burr, 2003): A classic constraint-based algorithm that identifies the causal graph’s skeleton using an independence test. (2) Dynotears (Pamfil et al., 2020): Constructs dynamic Bayesian networks through vector autoregression models. (3) C-LSTM (Tank et al., 2022): Utilizes LSTM to model temporal dependencies and capture nonlinear Granger causality. (4) GOLEM (Ng et al., 2020): Relaxes the hard Directed Acyclic Graph (DAG) constraint in NOTEARS (Zheng et al., 2018) with a scoring function. (5) REASON (Wang et al., 2023c): Learns both intra-level and inter-level causal relationships in interdependent networks. (6) MULAN (Zheng et al., 2024a): A multi-modal method that captures both modality-invariant and modality-specific representations. (7) CORAL (Wang et al., 2023a): A VAR-based online RCA method that decouples state-invariant and state-dependent information.
The first four baseline models were originally designed to learn causal structures solely from time series data. As outlined in (Wang et al., 2023c; a), these causal discovery models can be extended to identify the root cause nodes. In this process, we first apply causal discovery models to learn the causal graphs, then utilize random walk with restarts (Wang et al., 2023a) on these graphs to identify the top $K$ nodes as root causes.
Table 2: Results on Product Review dataset w.r.t different metrics.
@@ 4.1 Experimental Setup
GOLEM 0 0 0.25 0.043 0 0 0.025 308.25
REASON 0.25 1.0 1.0 0.5625 0.583 0.75 0.875 247.87
CORAL 0.5 1.0 1.0 0.75 0.833 0.9 0.95 146.46
@@
GOLEM 0 0 0.25 0.058 0 0 0.075 144.8
REASON 0 0 0.5 0.088 0 0 0.1 129.17
CORAL 0 0 0.5 0.118 0 0 0.2 50.29
@@
C-LSTM 0.25 0.5 0.5 0.341 0.25 0.35 0.425 1808.76
GOLEM 0 0 0.25 0.066 0 0 0.05 452.25
REASON 0.5 1.0 1.0 0.687 0.667 0.8 0.9 303.5
MULAN 0.75 1.0 1.0 0.833 0.833 0.9 0.95 255.74
CORAL 0.75 1.0 1.0 0.875 0.917 0.95 0.975 186.73
OCEAN 1.0 1.0 1.0 1.0 1.0 1.0 1.0 20.16
Table 3: Results on Online Boutique w.r.t different metrics.
@@
C-LSTM 0 0.4 0.8 0.3 0.1 0.2 0.44 20.75
GOLEM 0 0.4 0.6 0.291 0.2 0.267 0.36 4.32
REASON 0.4 0.8 1.0 0.617 0.5 0.6 0.76 3.23
CORAL 0.2 1.0 1.0 0.6 0.6 0.733 0.84 2.99
Log Only PC 0 0.4 0.6 0.257 0.1 0.2 0.32 3.88
@@
C-LSTM 0 0.4 0.6 0.267 0.1 0.2 0.36 15.07
GOLEM 0 0.4 0.8 0.248 0 0.133 0.36 3.39
REASON 0.2 0.8 0.8 0.458 0.3 0.467 0.6 2.39
CORAL 0.2 0.6 1.0 0.457 0.3 0.4 0.6 2.04
@@
MULAN 0.4 0.8 1.0 0.617 0.5 0.6 0.76 4.96
CORAL 0.4 1.0 1.0 0.7 0.7 0.8 0.88 3.63
OCEAN 0.6 1.0 1.0 0.8 0.8 0.867 0.92 1.84
4.2 Performance Evaluation
Experimental Results.
In this subsection, we present the performance evaluation results in Tables 2, 3, and 5 for various methods. Due to the page limit, we have moved the results for the Train Ticket dataset (i.e., Table 5) to Appendix E.1. Notably, although many baseline methods (e.g., PC, C-LSTM, REASON, Dynotears, GOLEM) are tailored for single-modal scenarios, we assess their performance in both single-modal contexts (e.g., system metrics only or system logs only) and multi-modal scenarios. System logs are considered additional system metrics, enabling these single-modal methods to be evaluated in a multi-modal context. We derive an average ranking score from different system metrics as the final result for all single-modal methods.
Our findings include: (1) Most baseline methods show improved performance when leveraging multi-modal data across three datasets. (2) CORAL, as an online RCA method, surpasses all offline methods across seven metrics. (3) OCEAN consistently outperforms all baselines across the datasets. (4) Both CORAL and OCEAN demonstrate shorter training time compared to offline methods, with OCEAN reducing its computational costs to 1/9 that of CORAL on the Product Review dataset. This reduction is credited to the efficiency of dilated convolutional operations and the design of the multi-factor attention module. CORAL’s approach of individually learning and then fusing causal graphs for each metric is computationally intensive in an online setting. Furthermore, OCEAN shows a notable improvement in MRR on the Product Review dataset, outperforming CORAL by 12.5%. Additionally, OCEAN exceeds CORAL by 20% in PR@1 and 10% in MAP@2 on the Online Boutique dataset, benefiting from the assessment of the importance of multiple factors and exploring correlations among different modalities.
Table 4: Ablation study on three datasets w.r.t MRR.
@@ 4.2 Performance Evaluation
Ablation Study.
In this subsection, we evaluate the effectiveness of individual components within the objective function of OCEAN (Eq. 3.4). Specifically, we define OCEAN-F and OCEAN-M as variants that lack the multi-factor attention learning module and the contrastive multi-modal learning module, respectively, while OCEAN-S removes the sparse constraint. The results, shown in Table 4, indicate a significant performance degradation when any component is omitted. Specifically, removing the multi-factor attention module results in 25% and 5% performance drop on the Product Review dataset and Train Ticket dataset, respectively. Eliminating the contrastive multi-modal learning module leads to 12.5% reduction on the Product Review dataset. These findings underscore the importance of each component in maintaining OCEAN’s high performance. Additionally, as detailed in Appendix E.2, replacing the dilated convolutional network with LSTM or Transformer model resulted in performance decline, reinforcing the effectiveness of our design.
Case Study.
@@
In this subsection, we evaluate the promptness of two online RCA methods on the Product Review dataset, CORAL and OCEAN, as shown in Figure 2 (a). Note that we also evaluate the effectiveness of the long-term temporal causal structure learning module by varying the number of dilated convolutional layers in OCEAN, specifically comparing configurations with one and two layers. In Figure 2 (a), the y-axis represents the batch index at which an RCA method meets the stopping criteria, and the real-time marker indicates the actual system failure time. A lower batch index value signifies faster identification of the ground-truth root cause by the RCA method. Notably, CORAL did not successfully rank the ground-truth root cause first in case 2, so we use the total number of batches to represent its detection time for a fair comparison. Our observations reveal that CORAL experiences about a 10-epoch delay relative to real-time in most cases, whereas OCEAN (2-layer) achieves quicker detection than OCEAN (1-layer). This improvement confirms our hypothesis that adding more dilated convolutional layers enhances the model’s ability to capture longer temporal dependencies, as discussed in Subsection 3.2.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Identification Time (b) $\ln(\lambda_{1})$ w.r.t. MRR (c) $\ln(\lambda_{2})$ w.r.t. MRR (d) $\ln(\lambda_{3})$ w.r.t. MRR
Figure 2: Figure (a) shows the identification time for four cases as well as the average identification time. Figures (b), (c), and (d) are the parameter analysis on the Product Review dataset w.r.t MRR.
Parameter Analysis.
In this subsection, we present detailed parameter sensitivity analysis conducted on the Product Review dataset, with additional analyses for other datasets available in Appendix E.3. Specifically, we explore the impact of three parameters, $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, on the overall objective functions as defined in Eq. 3.4. The experimental results are displayed in Figures 2 (b), (c), and (d), showing the Mean Reciprocal Rank (MRR) on the Product Review dataset. On these figures, the x-axis represents $\ln(\lambda_{i})$ for $i\in[1,2,3]$, and the y-axis shows the MRR score.
Our analysis reveals that a higher $\lambda_{1}$ (e.g., $\ln(\lambda_{1})=4.6$) significantly enhances performance, underscoring the vital role of the long-term temporal causal structure learning module in capturing temporal dependencies among system entities. Conversely, $\lambda_{2}$ and $\lambda_{3}$ exhibit optimal performance at relatively lower values (e.g., $\ln({\lambda_{2}})=-0.7$ and $\ln(\lambda_{3})=0$), with performance declining noticeably at higher levels. However, further reducing $\lambda_{2}$ and $\lambda_{3}$ also leads to diminished performance, which verifies the important role of sparse regularization and the acyclic constraint of the causal graph.
5 Conclusion
@@
Appendix A More Rational Behind the Model Design
The main issue of Recurrent Neural Networks (RNNs) and Transformers is their high computational cost, which makes them impractical for online root cause analysis. We chose Dilated Convolutional Neural Networks (DCNNs) for their efficiency and ability to capture long-range dependencies, both of which are essential in online settings where computational efficiency is critical. Unlike RNNs and Transformers, which could capture temporal dependencies but often come with high computational overhead, DCNNs are parallelizable and offer a lower time complexity of $O(NTkC)$, where $T$ is the sequence length, $L$ is the number of layers, $C$ is the number of filters, and $k$ is the filter size. In contrast, the complexity of the Transformer is $O(LT^{2}d)$, where $T$ is the sequence length, $L$ is the number of layers, and $d$ is the hidden feature dimensionality. As a result, Transformers are computationally more expensive than dilated convolutional neural networks. We further demonstrate the efficiency and effectiveness of our method compared to LSTM and Transformers in the ablation study presented in Table 6.
The rational behind the multi-factor attention mechanism is to capture complex relationships among various factors (e.g., CPU usage and memory usage metrics in a microservice system) across two modalities. A common limitation of existing methods is their focus on correlations between modalities, often overlooking the importance of individual factors within each modality. Our mechanism addresses this by reassessing the significance of each factor, enhancing causal structure learning. This enables our model to dynamically adjust to the changing significance of factors, which is crucial for accurate root cause analysis.
We integrate contrastive learning into our model because of its proven effectiveness in multi-modal tasks. It extracts shared information across different modalities, enhancing robustness and addressing challenges posed by low-quality data. By assigning weights to each modality based on its importance, contrastive learning ensures that critical patterns are not obscured by noise. This approach helps maintain the quality of causal graph learning, even when dealing with data of varying quality.
Appendix B Time Complexity
The time complexity of dilated convolution based causal structure learning is $O(NTkC)$. Here, $T$ is the sequence length, $L$ is the number of layers, $C$ is the number of filters, and $k$ is the filter size. The time complexity of multi-factor attention module is $O(d_{M}d_{L}d)$, where $d$ is the hidden feature dimensionality, and $d_{M}$ and $d_{L}$ are the number of factors in two modalities. The time complexity of contrastive learning module is $O(n^{2}d)$, where $n$ is the number of system entities.
Appendix C Log Feature/Indicator Extraction
In this subsection, we provide the details of converting the raw data into the time series, though this is not within the scope of this work. Specifically, we first use the Drain parser He et al. (2017) to transform the unstructured log event into structured log templates for each entity. Then, we partition the log data with fixed time windows, such as 5 minutes, and set time steps at 10 seconds. Within these intervals, we count the occurrence of each log template to derive the log frequency feature. The extraction of the log frequency feature is inspired by the insight that the recurrence of a log event template often correlates with its significance. For instance, when a microservice system experiences Distributed Denial of Service (DDoS) attacks, the system will produce an unusual volume of system logs, indicating abnormal activity. Thus, the log frequency provides the information to identify unusual patterns indicative of potential failure scenarios. In addition to log frequency, we also extract a second type of log feature known as the ‘golden signal.’ Notice that different from log frequency, golden signal heavily relies on domain knowledge and it only focuses on the abnormal system logs. More specifically, we are only interested in some keywords, including ‘error,’ ‘exception,’ ‘critical,’ ‘fatal’, and various others indicative of system anomalies. By identifying these keywords within log event templates, we can discern abnormal occurrences for system failure localization. Similar to the frequency-based feature, we compute the number of abnormal log events to derive the golden signal-based feature.
Appendix D Evaluation Metrics

### v2

@@ III-D Network Propagation-based Root Cause Identification
Malfunction effects often propagate from the root cause to its neighboring entities, implying that the immediate neighbors of anomalous KPIs may not necessarily be the true root causes. To accurately identify the root cause, we first derive a transition probability matrix based on the learned causal graph $\mathcal{G}$ and then employ a random walk with restart (RWR) algorithm [DBLP:conf/icdm/TongFP06] to simulate the propagation of malfunction signals:
$\mathbf{P}_{ij}=\frac{\mathcal{A}_{j,i}}{\sum_{k=1}^{n}\mathcal{A}_{k,i}},$ (18)
where $\mathbf{P}$ is the normalized adjacency matrix. During the random walk process, the model periodically restarts from the KPI node with probability $c\in[0,1]$ to explore alternative propagation paths. The RWR update equation is formulated as:
$\mathbf{r}_{t+1}=(1-c)\mathbf{P}\mathbf{r}_{t}+c\mathbf{r}_{0},$ (19)
where $\mathbf{r}_{t}$ denotes the probability distribution over nodes at step $t$, and $\mathbf{r}_{0}$ is the initial distribution concentrated on the target KPI node. Upon convergence, the stationary distribution $\mathbf{r}_{t+1}$ reflects the influence score of each entity, and the top-$k$ entities with the highest scores are identified as the most probable root causes.
Stopping Criterion. As the online RCA process progresses across multiple data batches, both the inferred causal structure and the corresponding root cause rankings tend to stabilize. To avoid redundant computation, we introduce an adaptive stopping mechanism based on the rank-biased overlap (RBO) metric [DBLP:journals/tois/WebberMZ10], which quantifies the similarity between two ranked lists while emphasizing higher-ranked items. Given the root cause lists from consecutive batches, $R_{t-1}$ and $R_{t}$, their similarity is computed as:
$\gamma$ $=\text{RBO}(R_{t-1},R_{t}),$ (20)
@@
where $\gamma\in[0,1]$. A larger $\gamma$ indicates a higher degree of stability in the root cause rankings. The online RCA process terminates automatically when the similarity score $\gamma$ exceeds a predefined threshold.
IV Experiments
@@
IV-A Experimental Setup
TABLE II: Results on Product Review dataset w.r.t different metrics.
@@ IV-A Experimental Setup
REASON 0.25 1.00 1.00 0.563 0.583 0.75 0.875 247.87
CORAL 0.50 1.00 1.00 0.750 0.833 0.90 0.950 146.46
Baro 0.50 0.50 0.75 0.537 0.500 0.50 0.550 283.48
$\epsilon$-Diagnosis 0 0 0.25 0.038 0 0 0.050 367.13
@@
REASON 0 0 0.5 0.088 0 0 0.100 129.17
CORAL 0 0 0.5 0.118 0 0 0.200 50.29
Baro 0.25 0.25 0.25 0.286 0.250 0.25 0.250 138.94
$\epsilon$-Diagnosis 0 0 0.25 0.039 0 0 0.050 149.23
@@
C-LSTM 0.25 0.5 0.5 0.341 0.250 0.35 0.425 1808.76
REASON 0.5 1.00 1.00 0.687 0.667 0.80 0.900 303.5
MULAN 0.75 1.00 1.00 0.833 0.833 0.90 0.950 255.74
CORAL 0.75 1.00 1.00 0.875 0.917 0.95 0.975 186.73
Baro 0.50 0.75 1.00 0.587 0.500 0.60 0.700 307.26
$\epsilon$-Diagnosis 0 0 0.25 0.042 0 0 0.075 402.33
OCEAN 1.00 1.00 1.00 1.000 1.000 1.00 1.000 20.16
TABLE III: Results on Online Boutique w.r.t different metrics.
@@
C-LSTM 0 0.4 0.8 0.300 0.1 0.200 0.44 20.75
REASON 0.4 0.8 1.0 0.617 0.5 0.600 0.76 3.23
CORAL 0.2 1.0 1.0 0.600 0.6 0.733 0.84 2.99
Baro 0 0.8 1.0 0.417 0 0.467 0.68 3.46
$\epsilon$-Diagnosis 0 0.6 1.0 0.323 0 0.267 0.52 5.21
Log Only PC 0 0.4 0.6 0.257 0.1 0.200 0.32 3.88
@@
C-LSTM 0 0.4 0.6 0.267 0.1 0.200 0.36 15.07
REASON 0.2 0.8 0.8 0.458 0.3 0.467 0.60 2.39
CORAL 0.2 0.6 1.0 0.457 0.3 0.400 0.60 2.04
Baro 0 0.6 0.8 0.308 0 0.267 0.48 2.57
$\epsilon$-Diagnosis 0 0.4 0.4 0.208 0 0.133 0.24 3.47
@@
MULAN 0.4 0.8 1.0 0.617 0.5 0.600 0.76 4.96
CORAL 0.4 1.0 1.0 0.700 0.7 0.800 0.88 3.63
Baro 0.2 1.0 1.0 0.567 0.2 0.667 0.80 4.88
$\epsilon$-Diagnosis 0 0.8 1.0 0.383 0 0.400 0.64 6.51
OCEAN 0.6 1.0 1.0 0.800 0.8 0.867 0.92 1.84
Datasets. We evaluate the performance of OCEAN using three public real-world datasets: (1) Product Review***https://lemma-rca.github.io/docs/data.html [zheng2024lemma]: A microservice system dedicated to online product reviews, encompassing $234$ pods deployed across $6$ cloud servers. It recorded four system faults between May 2021 and December 2021. (2) Online Boutiquehttps://github.com/IntelligentDDS/Nezha [yu2023nezha]: A microservice system designed for e-commerce, including five system faults. (3) Train Ticket [yu2023nezha]: A microservice system for railway ticketing services, also with five system faults. All three datasets contain two modalities: system metrics and system logs.
Evaluation Metrics. We evaluate the effectiveness of the proposed model with three widely-used metrics [wang2023interdependent, DBLP:conf/iwqos/MengZSZHZJWP20]: (1) Precision@K (PR@K): This metric measures the probability that the top-K predicted root causes are accurate. (2) Mean Average Precision@K (MAP@K): It assesses the top-K predicted causes from an overall perspective. (3) Mean Reciprocal Rank (MRR): This metric evaluates the ranking capability of the models. (4) Time: the training time (in seconds) for each batch of data.
Baselines. We compare OCEAN with eight methods: single modality RCA methods, including Dynotears [DBLP:conf/aistats/PamfilSDPGBA20],C-LSTM [DBLP:journals/pami/TankCFSF22], REASON [wang2023interdependent], multi-modality RCA method, i.e., MULAN [DBLP:conf/www/ZhengCHC24], hypothesis based methods, PC [DBLP:journals/technometrics/Burr03], $\epsilon$-Diagnosis [shan2019diagnosis], BARO [pham2024baro], one online RCA method (i.e., CORAL [wang2023incremental]).
Reproducibility. All experiments are conducted on a desktop running Ubuntu 18.04.5 with an Intel(R) Xeon(R) Silver 4110 CPU @2.10GHz and one 11GB GTX2080 GPU. We use the Adam as the optimizer and we train the model for 100 iterations at each batch. We use two layers of dilated convolutional operations in the experiment. As for the stopping criteria, we terminate the identification process if the similarity $\gamma$ between the current batch and the previous batch is greater than 0.9 for three consecutive times. Each experiment is run with one trial. For Product Review dataset, we set the size of historical metric and log data to 8-hour intervals and each batch is set to be a 10-minute interval. Due to page limitation, we include the additional experiments, such as the experiment on Train Ticket dataset, the comparison with physical graph, time complexity analysis, three-modalities analysis, comparison with different architecture, and parameter analysis in a more complete version onlinehttps://arxiv.org/abs/2410.10021.
IV-B Performance Evaluation
Experimental Results
In this subsection, we present the performance evaluation on Tables II and III for various methods. Considering that many baseline methods (e.g., PC, C-LSTM, REASON, Dynotears, and Baro) are designed for the single-modality scenario, we assess their performance in both single-modality scenarios (e.g., system metrics only or system logs only) and the multi-modality case. We consider system logs as additional system metrics to enable the performance measurement of these single-modality RCA methods in the multi-modality scenario. We calculate an average ranking score based on the evaluation of different system metrics as the final result for all single-modality methods and OCEAN. Our observations are as follows: (1) Compared to single-modality scenarios, most baseline methods benefit from leveraging multi-modality data across three distinct datasets. (2) As an online RCA method, CORAL outperforms all of the offline RCA methods with respect to seven metrics. (3). OCEAN consistently outperforms all baseline methods across the three datasets. (4). The online RCA methods CORAL and OCEAN have less training time compared with the offline RCA methods, while OCEAN further reduces its computational cost to 1/9 compared with CORAL. The reduced computational cost is attributed to the efficiency of dilated convolutional operation and the design of the multi-factor attention mechanism. Notice that CORAL first learns the causal graph for each factor individually and then fuses the causal graphs, which is computationally expensive in the online setting. Notably, OCEAN exhibits a remarkable improvement in MRR on the Product Review dataset, excelling the second competitor (i.e., CORAL) by 12.5%. Moreover, OCEAN outperforms CORAL by 20% and 10% with respect to PR@1 and MAP@3, respectively. This is attributed to the assessment of multiple factors and the exploration of the correlation among different modalities.
Figure 2: Identification time for four cases as well as the average identification time.
TABLE IV: Ablation study on three datasets w.r.t MRR.
@@ IV-B Performance Evaluation
Case Study
@@
In this subsection, we evaluate the promptness of two online RCA methods on the Product Review dataset, CORAL and OCEAN, as shown in Figure 2. Note that we also evaluate the effectiveness of the long-term temporal causal structure learning module by varying the number of dilated convolutional layers in OCEAN, specifically comparing configurations with one and two layers. In Figure 2, the y-axis represents the batch index at which an RCA method meets the stopping criteria, and the real-time marker indicates the actual system failure time. A lower batch index value signifies faster identification of the ground-truth root cause by the RCA method. Notably, CORAL did not successfully rank the ground-truth root cause first in case 2, so we use the total number of batches to represent its detection time for a fair comparison. Our observations reveal that CORAL experiences about a 10-epoch delay relative to real-time in most cases, whereas OCEAN (2-layer) achieves quicker detection than OCEAN (1-layer). This improvement confirms our hypothesis that adding more dilated convolutional layers enhances the model’s ability to capture longer temporal dependencies, as discussed in Subsection III-B.
Ablation Study
In this subsection, we evaluate the effectiveness of individual components within the objective function of OCEAN (Eq. III-C). Specifically, we define OCEAN-F and OCEAN-M as variants that lack the multi-factor attention learning module and the contrastive multi-modal learning module, respectively, while OCEAN-S removes the sparse constraint. The results, shown in Table IV, indicate a significant performance degradation when any component is omitted. Specifically, removing the multi-factor attention module results in 25% and 5% performance drop on the Product Review dataset and Train Ticket dataset, respectively. Eliminating the contrastive multi-modal learning module leads to 12.5% reduction on the Product Review dataset. These findings underscore the importance of each component in maintaining OCEAN’s high performance.
V Conclusion
@@
Appendix A Appendix
TABLE V: Results on Train Ticket w.r.t different metrics.
Modality Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10 Time (s)
Metric Only PC 0 0 0.2 0.067 0 0 0.060 9.65
Dynotears 0 0 0 0.047 0 0 0 21.3
C-LSTM 0 0.2 0.2 0.097 0 0.08 0.140 30.63
REASON 0.2 0.4 0.6 0.323 0.2 0.36 0.480 9.03
CORAL 0 0.4 1.0 0.184 0 0.16 0.500 5.72
Baro 0 0.2 0.8 0.145 0 0.08 0.243 9.68
$\epsilon$-Diagnosis 0 0.2 0.4 0.105 0 0.08 0.143 11.32
Log Only PC 0 0.2 0.4 0.166 0.13 0.16 0.260 6.34
Dynotears 0 0 0.2 0.072 0 0 0.020 17.26
C-LSTM 0 0 0.2 0.072 0 0 0.020 21.40
REASON 0 0.2 0.6 0.126 0 0.08 0.280 7.58
CORAL 0 0.2 0.8 0.138 0 0.08 0.320 3.35
Baro 0 0.4 0.8 0.175 0 0.16 0.314 7.89
$\epsilon$-Diagnosis 0 0.2 0.4 0.106 0 0.08 0.114 9.77
Multi-Modality PC 0 0 0.2 0.083 0 0 0.100 12.83
Dynotears 0 0.4 0.6 0.141 0 0.16 0.320 27.82
C-LSTM 0.2 0.4 0.6 0.294 0.2 0.28 0.360 36.76
REASON 0.2 0.4 0.6 0.300 0.2 0.28 0.420 12.81
MULAN 0.2 0.4 1.0 0.317 0.2 0.28 0.460 11.42
CORAL 0.2 0.4 1.0 0.334 0.2 0.28 0.560 7.26
Baro 0 0.4 1.0 0.254 0.33 0.36 0.414 13.74
$\epsilon$-Diagnosis 0 0 0.6 0.114 0 0 0.143 16.28
OCEAN 0.2 0.4 1.0 0.381 0.33 0.36 0.580 3.22
TABLE VI: Results on Train Ticket w.r.t different metrics.
Modality Model PR@1 PR@5 PR@10 MRR MAP@3 MAP@5 MAP@10 Time (s)
Metric Only PC 0 0 0.2 0.067 0 0 0.060 9.65
Dynotears 0 0 0 0.047 0 0 0 21.3
C-LSTM 0 0.2 0.2 0.097 0 0.08 0.140 30.63
REASON 0.2 0.4 0.6 0.323 0.2 0.36 0.480 9.03
CORAL 0 0.4 1.0 0.184 0 0.16 0.500 5.72
Baro 0 0.2 0.8 0.145 0 0.08 0.243 9.68
$\epsilon$-Diagnosis 0 0.2 0.4 0.105 0 0.08 0.143 11.32
Log Only PC 0 0.2 0.4 0.166 0.13 0.16 0.260 6.34
Dynotears 0 0 0.2 0.072 0 0 0.020 17.26
C-LSTM 0 0 0.2 0.072 0 0 0.020 21.40
REASON 0 0.2 0.6 0.126 0 0.08 0.280 7.58
CORAL 0 0.2 0.8 0.138 0 0.08 0.320 3.35
Baro 0 0.4 0.8 0.175 0 0.16 0.314 7.89
$\epsilon$-Diagnosis 0 0.2 0.4 0.106 0 0.08 0.114 9.77
Multi-Modality PC 0 0 0.2 0.083 0 0 0.100 12.83
Dynotears 0 0.4 0.6 0.141 0 0.16 0.320 27.82
C-LSTM 0.2 0.4 0.6 0.294 0.2 0.28 0.360 36.76
REASON 0.2 0.4 0.6 0.300 0.2 0.28 0.420 12.81
MULAN 0.2 0.4 1.0 0.317 0.2 0.28 0.460 11.42
CORAL 0.2 0.4 1.0 0.334 0.2 0.28 0.560 7.26
Baro 0 0.4 1.0 0.254 0.33 0.36 0.414 13.74
$\epsilon$-Diagnosis 0 0 0.6 0.114 0 0 0.143 16.28
OCEAN 0.2 0.4 1.0 0.381 0.33 0.36 0.580 3.22
A-A Evaluation Metrics